OpenStudy (anonymous):

when sin(x) + cos(x)= (1/2) What is sin 2x

OpenStudy (anonymous):

Use this sine property: \[\sin(2x) = 2\sin(x)\cos(x)\] First, isolate either sine or cosine from the given equation then sub it into this sine property.